# A Calculus of Inheritance

### 10 Mar 2026 v6

@@ 4. Embedding the $\lambda$-Calculus
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results (and, optionally, naming values), and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
@@
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the application $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the value $\mathbf{let}$-binding rule, $x$ wraps $\mathcal{T}(V)$ under a $\mathrm{result}$ child so that the uniform accessor $x.\mathrm{result}$ yields $\mathcal{T}(V)$, the translated value itself. This extra wrapper is semantically inert but proof-relevant: it lets the same $\mathrm{result}$-following argument used for applications extend uniformly to value $\mathbf{let}$-bindings in the adequacy proof below. In the tail-call rule, the fresh label $\mathrm{tailCall}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In the application and tail-call cases, the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The six rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus.
@@ 4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces an elementary semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
A mixin tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
Definition 4.1 (Inheritance-convergence).
in the semantics of Section 3.
Theorem 4.2 (Adequacy).
Theorem 4.3 (First-order semantics of the lazy $\lambda$-calculus).
For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ iff for every closing $\lambda$-calculus context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$. Then:
Here $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2). The theorem establishes that the lazy $\lambda$-calculus has a first-order semantics: the translation $\mathcal{T}$ embeds the $\lambda$-calculus into a sublanguage of inheritance-calculus, using only recursive evaluation of set-comprehension equations over paths and sets of paths. The prior higher-order semantic frameworks of (Scott, 1972; Wadsworth, 1976; Abramsky et al., 2000; Hyland and Ong, 2000; Barendregt, 1984) are thereby shown to be faithful to this first-order semantics: they agree with the equations on Böhm tree equivalence (Barendregt, 1984). [^7] Composing $\mathcal{T}$ with the equations of Section 3 yields a first-order semantics of the $\lambda$-calculus in which function application is record inheritance and convergence is determined by path queries. Appendix E reformulates this semantics directly on ANF, without passing through inheritance-calculus. The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix B.
@@ 4.2. Expressive Asymmetry
The underlying mechanism is open extension: mixin definitions compose via open recursion, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed: there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
@@ 5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026). This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names serve as data constructors and operations; lowerCamelCase names serve as fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
@@
Following the Church-encoding convention (Yang, 2026), [^8] The data representation (Zero/Successor) is a Peano encoding, an initial algebra of the functor $F(X)=1+X$. The visitor pattern (NatVisitor, below) adds the Church-encoding behavior: a fold (catamorphism) over this algebra. These two concerns are defined in separate files and composed via inheritance, which is itself an instance of the Expression Problem. we define a factory containing the abstract product type and two labels that serve as data constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The file name $\mathrm{NatData}$ is itself the top-level scope:
@@
Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each data constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
NatEquality
To test that $2+3=5$, we needed equality on Nat values. Equality requires case analysis on the data constructor; we use the visitor pattern. $\mathrm{NatVisitor}$ adds an $\mathrm{Acceptance}$ operation to each data constructor, containing a $\mathrm{VisitorMap}$ with the appropriate $\mathrm{ZeroVisitor}$ or $\mathrm{SuccessorVisitor}$ branch, and an $\mathrm{Accepted}$ field that delegates to the matching branch:
@@
$\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each data constructor. $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor. Case analysis on $\mathrm{other}$ is done via its $\mathrm{Acceptance}$:
@@
Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either file, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically. The concrete numerals live in a shared file that inherits only $\mathrm{NatData}$:
@@
Querying $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ via the host-language API returns $\mathrm{True}$. No existing file was modified. $\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new file: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
@@
Each operation is an independent mixin that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every label in the factory acquires all operations from all inherited mixins. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-path qualified this (equation 6) is essential. In the test file above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$. When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes. The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix D), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-path inheritance to shared schemas, which DOT’s single-valued self variable cannot accommodate.
@@
A Church-encoded Nat is a record whose structure mirrors the data-constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on. This is exactly the shape of a trie. A scope that simultaneously inherits two Nat values is a trie union. The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$:
@@
$\mathrm{Result}.\mathrm{sum}$ evaluates to the trie $\{4,5,6\}$, the set of all pairwise sums. $\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; querying $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ via the host-language API returns $\{\mathrm{True},\mathrm{False}\}$. The Cartesian product arises directly from the semantic equations. $\mathrm{OneOrTwo}$ inherits both $\mathrm{One}$ and $\mathrm{Two}$, so $\mathtt{bases}(\mathrm{OneOrTwo})$ contains two paths. When $\mathrm{Result}$ inherits $\mathrm{OneOrTwo}.\mathrm{Plus}$, the $\mathtt{bases}$ of $\mathrm{Result}$ include the $\mathrm{Plus}$ of both $\mathrm{One}$ and $\mathrm{Two}$. Each $\mathrm{Plus}$ resolves its $\mathrm{addend}$ via $\mathtt{this}$, reaching $\mathrm{ThreeOrFour}$, which is itself a trie union of $\mathrm{Three}$ and $\mathrm{Four}$: the $\mathtt{supers}$ of the $\mathrm{addend}$ path contain both values. Since $\mathrm{Successor}.\mathrm{Plus}$ recursively delegates to $\mathrm{predecessor}.\mathrm{Plus}$ with an incremented addend, and $\mathrm{predecessor}$ also resolves through the trie union, the recursion branches at every trie node. The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overrides}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.

### 16 Mar 2026 v7

@@ 4. Embedding the $\lambda$-Calculus
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results. [^8] This can optionally include naming values as well; since every ANF term is also a $\lambda$-term, ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
@@
An abstraction $\lambda x.\,M$ translates to a record with own properties $\{\mathrm{argument},\mathrm{result}\}$; we call this the abstraction shape. In the application $\mathbf{let}$-binding rule, the name $x$ binds to the inherited record $\{\mathcal{T}(V_{1}),\;\mathrm{argument}\mapsto\mathcal{T}(V_{2})\}$, and $\mathcal{T}(M)$ may reference $x.\mathrm{result}$ to obtain the value of the application. In the value $\mathbf{let}$-binding rule, $x$ wraps $\mathcal{T}(V)$ under a $\mathrm{result}$ child so that the uniform accessor $x.\mathrm{result}$ yields $\mathcal{T}(V)$, the translated value itself. This extra wrapper is semantically inert but proof-relevant: it lets the same $\mathrm{result}$-following argument used for applications extend uniformly to value $\mathbf{let}$-bindings in the adequacy proof (Appendix B). In the tail-call rule, the fresh label $\mathrm{tailCall}$ serves the same purpose, and $\mathrm{result}$ projects the answer. In the application and tail-call cases, the inheritance that performs the application is encapsulated behind a named property, so its internal structure (the $\mathrm{argument}$ and $\mathrm{result}$ labels of the callee) does not leak to the enclosing scope.
The six rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus. [^10] MIXINv2 separates symbol resolution (which resolves lexical references and qualified $\mathtt{this}$ references) from runtime mixin evaluation, but symbol resolution is performed just-in-time per symbol rather than ahead of time, so it does not provide static type checking in the traditional sense. Nevertheless, MIXINv2’s resolution phase is stricter than the untyped inheritance-calculus. All examples in this paper are ported from MIXINv2 and conform to both semantics. Elements added to satisfy the resolver are harmless for the algorithms presented here and may aid readability.
@@ 4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ embeds the $\lambda$-calculus into a sublanguage of inheritance-calculus. Intuitively, a mixin tree augments the AST with additional semantic structure; for translated $\lambda$-terms, the projection of this additional structure is isomorphic to the Böhm tree. A mixin tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. The translation preserves and reflects convergence (adequacy). Write $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ when $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$ for every closing $\lambda$-calculus context $C[\cdot]$; then $\approx_{\lambda}$ coincides with Böhm tree equivalence (Barendregt, 1984). Formal statements and proofs are in Appendix B. [^11] Appendix E applies the same first-order technique directly on ANF to give the $\lambda$-calculus a self-contained fixpoint semantics, without passing through inheritance-calculus. Note that $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2). Böhm tree equivalence is fully abstract for the first-iteration approximation $T_{P}\!\uparrow\!1$ of the equations. The full $\mathrm{lfp}(T_{P})$ is strictly more defined. For example, $\mathbf{let}\;x=x\;\mathbf{in}\;x$ diverges under $\beta$-reduction, but its translation $\{x\mapsto\{\mathrm{result}\mapsto x.\mathrm{result}\},\;\mathrm{result}\mapsto x.\mathrm{result}\}$ stabilizes under $\mathrm{lfp}(T_{P})$: [^12] This simple cycle could also be resolved by a fixed-point combinator such as $Y$. The $\mathrm{lfp}(T_{P})$ semantics is more general: it computes fixed points at the container level, e.g. finding a set $S$ such that $S=F(S)$. Recursive Datalog on cyclic graphs (Appendix F, Section F.8) requires exactly this, and neither $Y$ nor Haskell’s mfix (RecursiveDo) can express it. tabled evaluation detects the cycle and returns $\mathtt{properties}(\mathrm{root})=\{x,\mathrm{result}\}$ with no abstraction shape, reflecting an undefined value rather than diverging (Remark B.5). [^13] We call the sublanguage of inheritance-calculus corresponding to $\lambda$-terms, evaluated under $\mathrm{lfp}(T_{P})$ rather than $T_{P}\!\uparrow\!1$, the fixpoint $\lambda$-calculus; Remark B.5 elaborates.
@@ 4.2. Expressive Asymmetry
The underlying mechanism is open extension: mixin definitions compose via open extension, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed: there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
@@ 5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026). [^14] The examples in this section are rendered in inheritance-calculus notation; the executable MIXINv2 sources are available as supplementary material. This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one top-level mixin; UpperCamelCase names serve as data constructors and operations; lowerCamelCase names serve as fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
@@
The data representation (Zero/Successor) is a Peano encoding, an initial algebra of the functor $F(X)=1+X$. The visitor pattern (NatVisitor, below) adds Church-encoding behavior: a fold (catamorphism) over this algebra. These two concerns are defined in separate files and composed via inheritance (Yang, 2026), which is itself an instance of the Expression Problem. Following this convention, we define a factory containing the abstract product type and two labels that serve as data constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The name $\mathrm{NatData}$ is the top-level scope:
@@
Addition is a separate mixin that inherits $\mathrm{NatData}$ and extends each data constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
NatVisitor
Equality requires case analysis on the data constructor; we use the visitor pattern. $\mathrm{NatVisitor}$ adds an $\mathrm{Acceptance}$ operation to each data constructor, containing a $\mathrm{VisitorMap}$ with the appropriate $\mathrm{ZeroVisitor}$ or $\mathrm{SuccessorVisitor}$ branch, and an $\mathrm{Accepted}$ field that delegates to the matching branch:
@@
NatEquality
To test that $2+3=5$, we needed equality on Nat values. The Cartesian product test (page 5) will apply this same $\mathrm{Equal}$ to a trie union of multiple Nat values without modification. $\mathrm{NatEquality}$ inherits $\mathrm{NatVisitor}$ and $\mathrm{BooleanData}$, and adds an $\mathrm{Equal}$ operation to each data constructor. $\mathrm{Zero}$ is equal only to another $\mathrm{Zero}$; $\mathrm{Successor}$ is equal only to another $\mathrm{Successor}$ with an equal predecessor. Case analysis on $\mathrm{other}$ is done via its $\mathrm{Acceptance}$:
@@
Composing $\mathrm{NatPlus}$ with $\mathrm{NatEquality}$ by inheriting both, with no modifications to either mixin, gives the resulting Nat values both $\mathrm{Plus}$ and $\mathrm{Equal}$ automatically. The concrete numerals live in a shared mixin that inherits only $\mathrm{NatData}$:
@@
The path $\mathrm{Test}.\mathrm{Test2plus3}.\mathrm{equal}$ inherits $\mathrm{True}$. No existing mixin was modified. $\mathrm{NatEquality}$, a new operation, and $\mathrm{Boolean}$, a new data type, each required only a new mixin: extensibility in both dimensions of the Expression Problem (Wadler, 1998).
@@
Each operation is an independent mixin that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every label in the factory acquires all operations from all inherited mixins. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either mixin. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
Multi-target qualified $\mathtt{this}$ (equation 6) is essential. In the test mixin above, $\mathrm{NatPlus}.\mathrm{Successor}$ and $\mathrm{NatEquality}.\mathrm{Successor}$ both inherit $\mathrm{NatData}.\mathrm{Successor}$. When the two operations are composed, the resulting $\mathrm{Successor}$ inherits the $\mathrm{NatData}$ schema through two independent routes. The reference $\mathrm{Successor}.\mathbf{this}.\mathrm{predecessor}$ inside each operation resolves through both routes, and since inheritance is idempotent, all routes contribute the same properties. In Scala, the same pattern, where two independently defined inner classes extend the same outer class’s trait, triggers a “conflicting base types” rejection (Appendix D), because DOT’s self variable is single-valued and cannot resolve through multiple inheritance routes simultaneously. The Expression Problem can still be solved in Scala using object algebras or similar patterns, but composing independently defined operations inherently produces multi-target $\mathtt{this}$ resolution to shared schemas, which DOT’s single-valued self variable cannot accommodate.
@@
A Church-encoded Nat is a record whose structure mirrors the data-constructor path used to build it: $\mathrm{Zero}$ is a flat record; $\mathrm{S}(\mathrm{Zero})$ is a record with a $\mathrm{predecessor}$ field pointing to a $\mathrm{Zero}$ record; and so on. This is exactly the shape of a trie. [^15] The unary Nat trie can be viewed as a sparse linked list whose depth equals the number it represents: in a trie union, some $\mathrm{Successor}$ nodes along the list point to $\mathrm{Zero}$ (leaf nodes), so not every position in the list is occupied. The MIXINv2 source (Yang, 2026) includes a binary natural number (BinNat) arithmetic whose trie depth is logarithmic. A scope that simultaneously inherits two Nat values is a trie union. The following test applies $\mathrm{Plus}$ to a trie union of $\{1,2\}$ with addend a trie union of $\{3,4\}$, then compares the result to itself with $\mathrm{Equal}$:
@@
$\mathrm{Result}.\mathrm{sum}$ evaluates to the trie $\{4,5,6\}$, the set of all pairwise sums. $\mathrm{Equal}$ then applies to every pair drawn from $\{4,5,6\}\times\{4,5,6\}$; the path $\mathrm{CartesianTest}.\mathrm{Check}.\mathrm{equal}$ inherits both $\mathrm{True}$ and $\mathrm{False}$. The Cartesian product arises directly from the semantic equations. $\mathrm{OneOrTwo}$ inherits both $\mathrm{One}$ and $\mathrm{Two}$, so $\mathtt{bases}(\mathrm{OneOrTwo})$ contains two paths. When $\mathrm{Result}$ inherits $\mathrm{OneOrTwo}.\mathrm{Plus}$, the $\mathtt{bases}$ of $\mathrm{Result}$ include the $\mathrm{Plus}$ of both $\mathrm{One}$ and $\mathrm{Two}$. Each $\mathrm{Plus}$ resolves its $\mathrm{addend}$ via $\mathtt{this}$, reaching $\mathrm{ThreeOrFour}$, which is itself a trie union of $\mathrm{Three}$ and $\mathrm{Four}$: the $\mathtt{supers}$ of the $\mathrm{addend}$ path contain both values. Since $\mathrm{Successor}.\mathrm{Plus}$ recursively delegates to $\mathrm{predecessor}.\mathrm{Plus}$ with an incremented addend, and $\mathrm{predecessor}$ also resolves through the trie union, the recursion branches at every trie node. The result $\mathrm{sum}$ therefore collects all paths reachable through any combination of the two operands, which is the Cartesian product.
The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: [^16] Appendix F gives the systematic encoding of Datalog into inheritance-calculus. $\mathtt{overrides}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.